Solve for x
2(x-4)-(x+2)=(x+2)(x-4)

use a comma to separate answers as needed

Answer by mananth(16949)

2(x-4)-(x+2)=(x+2)(x-4)
...
2x-8-x-2=x^2-2x-8
x^2-3x+2=0
x^2-2x-x+2=0
x(x-2)-1(x-2)=0
(x-2)(x-1)=0
x = 1, 2
...
m.ananth@hotmail.ca